The differential equation satisfied by a beam uniformly loaded ( W kg/ meter ) with one end fixed and the second end subjected to a tensile force-P is given by
El d^2 y/dx^2 = Py - 1/2 Wx^2
Where E is the modulus of elasticity and l is the moment of inertia show that the elastic curve for the beam with condition y = 0 and dy/dx = 0 at x = 0 is given by
y = W/Pn^2 (1-cosh nx ) + Wx^2 /2P where n^2 = ( P/EI)
1
Expert's answer
2020-03-11T14:16:31-0400

Transform the original equation to get less confused with constants



"\left.EL\cdot\frac{d^2y}{dx^2}=P\cdot y-\frac{Wx^2}{2}\right|\times\frac{1}{EL}\longrightarrow\\[0.3cm]
y''=\overbrace\frac{P}{EL}^{n^2}\cdot y-\overbrace\frac{P}{EL}^{n^2}\cdot\frac{Wx^2}{2P}\longrightarrow\\[0.3cm]
\boxed{y''=n^2y-\frac{Wn^2x^2}{2P}}"



1 STEP: Solve the homogeneous equation

"y''=n^2y\to[y(x)=e^{kx}\,\,\,\text{and}\,\,\,y''=k^2\cdot e^{kx}]\to\\[0.3cm]
k^2\cdot e^{kx}=n^2\cdot e^{kx}\to\left[\begin{array}{c}
 k_1=n\\
 k_2=-n
\end{array}\right."



Then,


"\boxed{y_{hom}(x)=C_1\cdot e^{nx}+C_2\cdot e^{-nx}}"



2 STEP: Solve an inhomogeneous equation

To do this, find ANY particular solution, so we will look for it in the form



"y_{par}(x)=Ax^2+Bx+C\to y''=2A"



Substitute everything in the original equation



"2A=n^2\cdot\left(Ax^2+Bx+C\right)-\frac{Wn^2x^2}{2P}\to\\[0.3cm]
\left(An^2-\frac{Wn^2}{2P}\right)x^2+Bn^2x+(Cn^2-2A)=0\to\\[0.3cm]
\left\{\begin{array}{l}
An^2-\displaystyle\frac{Wn^2}{2P}=0\to\boxed{A=\frac{W}{2P}}\\[0.3cm]
Bn^2=0\to\boxed{B=0}\\[0.3cm]
Cn^2-2A=0\to\boxed{C=\frac{W}{Pn^2}}
\end{array}\right."

Conclusion,



"\boxed{y_{par}(x)=\frac{Wx^2}{2P}+\frac{W}{Pn^2}}\\[0.3cm]
y_{gen}(x)=y_{hom}(x)+y_{par}(x)\to\\[0.3cm]
\boxed{y_{gen}(x)=C_1\cdot e^{nx}+C_2\cdot e^{-nx}+\frac{Wx^2}{2P}+\frac{W}{Pn^2}}"

To find the constants we will use the boundary conditions



"\left\{\begin{array}{l}y(0)=C_1+C_2+\displaystyle\frac{W}{Pn^2}=0\\[0.3cm]
y'(0)=nC_1-nC_2=0\to C_1=C_2
\end{array}\right.\longrightarrow\\[0.3cm]
2C_1+\frac{W}{Pn^2}=0\to\boxed{C_1=C_2=-\frac{W}{2Pn^2}}"



Then,



"y(x)=-\frac{W}{2Pn^2}\cdot e^{nx}-\frac{W}{2Pn^2}\cdot e^{-nx}+\frac{Wx^2}{2P}+\frac{W}{Pn^2}\to\\[0.3cm]
y(x)=\frac{W}{2Pn^2}\cdot\left(2-\overbrace{\left(e^{nx}+e^{-nx}\right)}^{2\cosh nx}\right)+\frac{Wx^2}{2P}\to\\[0.3cm]
\boxed{y(x)=\frac{W}{Pn^2}\cdot(1-\cosh nx)+\frac{Wx^2}{2P}}"